Consumption function :C=200+0.9yd
Investment function :1=100
Government spending: G=250
Exports X=200
Imports 2= 50+0.12Y
Tax rate : T=20%
Calculate total autonomous expenditure
2.Hence using multiplier approach, calculate the equilibrium level of income Graphically illustrate it
3. Does government's budget exhibit a deficit or surplus? assume that the government collects income by means of taxes only
1
Expert's answer
2020-11-30T16:15:53-0500

1. Total autonomous expenditure

AE = C+I+G+NX

AE = 200+0.9yd+100+250+200

AE = 200+100+250+200+0.9yd

AE =750+0.9yd

2. Equilibrium level of income

Y= C+I+G

Y = 200+0.9yd+100+250

Y = 550+0.9yd

3. The government has a budget surplus because it collects more taxes than it spends.